OpenStudy (anonymous):

Multiply and simplify: square root of 6 times the quantity of square root of 3 + 5 square root of 2.

jimthompson5910 (jim_thompson5910):

The problem is this right? \[\Large \sqrt{6}*\left(\sqrt{3}+5\sqrt{2}\right)\]